OpenStudy (anonymous):

3y[1-(y-3)]-2[3-y(y-4)] simplfy

OpenStudy (anonymous):

Do the bits in ( )'s first and simplify there...

OpenStudy (anonymous):

how do you distribute the 1- to the (y-3) :( ??

OpenStudy (anonymous):

OK you you have 1 minus y minus minus 3 which is...?

OpenStudy (anonymous):

y+4?

OpenStudy (anonymous):

Good, now how about the other ( )?

OpenStudy (anonymous):

It's -y not y

OpenStudy (anonymous):

final answer : 5y^2 + 14y+24 ?

OpenStudy (anonymous):

you can just "read" it--- plus1 minusy minusminus3

OpenStudy (anonymous):

is my answer correct? im confused again

OpenStudy (anonymous):

You are going a bit too fast, do it more slowly and remember minus minus is a plus...

OpenStudy (anonymous):

so...your are saying that 3-y(y-4) = -y^2 -12 ?

OpenStudy (anonymous):

the y(y-4) you should first and then plus3

OpenStudy (anonymous):

Basically you want to get rid of those inside ( )'s before you do anything else, do you see what i mean?

OpenStudy (anonymous):

noo i dont understand.. y(y-4)+3 ?

OpenStudy (anonymous):

It doesn't say y(y-4) +3 it says -y(y-4) +3...

OpenStudy (anonymous):

Expand -y(y-4) and what do you get?

OpenStudy (anonymous):

-y^2-4

OpenStudy (anonymous):

-y^2-4y

OpenStudy (anonymous):

You seem to be getting stuck every time on minus minus, minus y times minus 4 equals PLUS 4y

OpenStudy (anonymous):

SOO 3y (y-4) -2 (-y^2+4y) :D ? nut there is the plus there for the second bracket go ? 3y (y-4) -2 (-y^2+4y)+3 ?

OpenStudy (anonymous):

Write the original question with the ( )'s expanded as we discussed and be careful with the signs

OpenStudy (anonymous):

3y^2-12y+2y^2-8y 5y^2-20y

OpenStudy (anonymous):

3y[ put something here ] -2 [ and here ]

OpenStudy (anonymous):

3y[ y-4] -2 [-y(y-4+3]

OpenStudy (anonymous):

i give up..

OpenStudy (anonymous):

Why do you put y-4?

OpenStudy (anonymous):

that what you said to write

OpenStudy (anonymous):

3y[ -y + 4 ] -2[ 3 - y^2 + 4y ] You need to understand this, then you can do the rest, OK

OpenStudy (anonymous):

thamks sexy\

OpenStudy (anonymous):

You're welcome:-)

OpenStudy (anonymous):

where u live

OpenStudy (anonymous):

final answer= y^2+4y-6

OpenStudy (anonymous):

????

OpenStudy (anonymous):

It's -y^2 not + y^2 (I live in Spain).

OpenStudy (anonymous):

ok ty